A rectangular field is three times as long as it is wide. if the perimeter of the field is 400 yards, what are the field`s dimentions?
Answer by xcentaur(357)

let length be x and width be y
then x=3y
2(x+y)=400
x+y=200
3y+y=200
4y=200
y=50
x=3*50=150


Dimensions are : 50 yards,150 yards


Hope this helps,
Prabhat